OpenStudy (anonymous):

14. Which equation is an identity? (1 point) 11 – (2v + 3) = –2v – 8 5w + 8 – w = 6w – 2(w – 4) 7m – 2 = 8m + 4 – m 8y + 9 = 8y – 3

OpenStudy (anonymous):

5w + 8 – w = 6w – 2(w – 4)

OpenStudy (anonymous):

The second one. An identity is always true. 5w + 8 - w =6w -2w +8 Now simplify and you get 4w +8 = 4w + 8 which is 4w=4w therefore it is always true.

OpenStudy (turingtest):

5w + 8 – w = 6w – 2(w – 4) 4w+8=6w-2w+8 4w+8=4w+8 0=0 which is always true, therefor this equation is an IDENTITY